# Spin Bott Indices For Time-Reversal-Invariant Higher-Order Topological Superconductors

### v1

@@ Abstract
The abundance of bulk and boundary topologies in higher-order topological phases offer remarkable tunability and diversity to boundary states but also pose a challenge to their unified topological characterization. In this work, we propose a theoretical framework to characterize time-reversal invariant topological superconductors hosting Majorana Kramers pairs (MKP) of corner states by using a series of spin Bott indices, which capture both bulk and boundary states topology. The developed invariants can characterize MKP in arbitrarily shaped systems and all distinct spatial distribution patterns of MKP. As an illustrative example, we apply our theory to analyze the Kane-Mele model with sublattice-dependent superconducting pairing potentials. In this representative model, various patterns of MKP can be engineered through edge cleavage, and despite their high sensitivity to boundary terminations, MKP can be faithfully characterized by the proposed topological invariants.
@@
Introduction.— Higher-order topological phases, hosting gapless boundary states with codimension greater than one, have attracted extensive research interest in both theory [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] and experiments [14, 15, 16, 17, 18, 19, 20]. Higher-order topological superconductors (HOTSCs) featuring Majorana corner states [21, 22, 23, 24, 25] are of particular interest because of their potential applications in topological quantum computation [26, 27, 28, 29]. Various systems have been proposed to realize HOTSCs, such as superconducting-proximitized topological insulators [30, 31, 32, 33, 34], odd-parity superconductors [35, 36, 37, 38], and iron-based superconductors [39, 40, 41, 42, 43, 44]. In contrast to first-order topology, the topological boundary states of HOTSCs typically depend on not only bulk states topology but also the boundary terminations [4]. The abundant bulk and boundary topologies of HOTSCs endow boundary states with high tunability and richness but also present the challenge of their unified topological characterizations [45, 46, 47, 48, 49, 50, 51, 52]. Several theoretical frameworks for characterizing higher-order topology have been developed, such as multipole moment [53, 54, 55], chiral multipole number [47], and symmetry indicators [56, 57, 58]. Despite these seminal works, a unified topological characterization of Majorana corner states remains elusive. A key difficulty is that the previously proposed invariants are defined by only bulk states while Majorana corner states can be sensitive to both bulk and boundary topologies [4, 59], which hinders establishing an exact correspondence between the proposed invariants and Majorana corner states.
In this Letter, we develop a theoretical framework to characterize time-reversal invariant HOTSCs by higher-order topological invariants based on a series of spin Bott indices. Our approach builds on our joint work [60], which proposed a method to characterize zero-energy corner states in systems with chiral symmetry by using Bott indices. Time-reversal invariant superconductors are in the DIII symmetry class and naturally host chiral symmetry $C=-iTP$, where $T$ and $P$ are, respectively, time-reversal and particle-hole symmetry. However, a direct application of the Bott index $N$ proposed in Ref. 60 based on the chiral symmetry $C$ of superconductors can not capture the higher-order topology since $N$ vanishes as required by the $T$ symmetry (see Eqs. (2) and (3)). Therefore, we develop spin Bott indices to characterize HOTSCs, in analogy to the spin-resolved topological characterization of quantum spin Hall insulators [61]. We emphasize that the Bott indices here are obtained under open boundary conditions, which enables capturing the topology of both bulk and boundary states. Particularly, the defined spin Bott indices can characterize MKP in systems with arbitrary shapes and all distinct spatial distribution patterns of MKP. To demonstrate the effectiveness of our theory, we study the Kane-Mele model [62] with sublattice-dependent superconducting pairing potentials. In this model, MKP can be flexibly engineered by boundary cleavage, as shown in Figs. 1 and 2, and all different patterns of MKP can be characterized by the proposed invariants. In addition, we also investigate disorder effects on the HOTSCs informed by the invariants.
@@
Spin Bott indices.— The symmetry operators of superconductors in the DIII symmetry class can be represented as $T=is_{y}\mathcal{K}$, $P=\tau_{y}s_{y}\mathcal{K}$, and $C=\tau_{y}$ with Pauli matrices $s_{y}$ and $\tau_{y}$ acting, respectively, on the spin and particle-hole spaces, and $\mathcal{K}$ being the complex conjugation operator. Recently, Bott indices have been proposed to characterize higher-order topological phases with chiral symmetry [47, 60]. We first show that the application of the previously defined Bott indices [47, 60] to superconductors in the DIII class based on $C$ can not capture the higher-order topology.
@@
Using singular value decomposition $h=U_{A}\Sigma U_{B}^{\dagger}$, the Bott index can be defined in the following equivalent ways [47, 60, 63]
@@
This restriction is consistent with the MKP related by $T$ having opposite eigenvalues of $C$. We denote the MKP at a given corner as $\gamma_{1}$ and $\gamma_{2}$, with $\gamma_{2}=T\gamma_{1}$. The Majorana corner states can always be chosen as the eigenstates of $C$ through a superposition of $\gamma_{1}$ and $\gamma_{2}$. Without loss of generality, we choose $C\gamma_{1}=\gamma_{1}$ and we have
@@
Here $M_{z}$ commutes with $T$. With $s_{z}$ conservation, $\mathcal{N}=\bar{\mathcal{N}}$ [63].
Characterization of MKP.— When $[s_{z},H]=0$, $H$ is block-diagonal in the spin space and each block belongs to the AIII symmetry class [64], allowing one corner to have $Z$ pairs of MKP. To describe the pattern of MKP at corners of a system with open-boundary conditions, we define the vector
@@
${f}_{4}^{(1)}=2\tilde{x}\tilde{y}/L,\quad{f}_{4}^{(2)}=\tilde{x}/L,\quad{f}_{4}^{(3)}=\tilde{y}/L,$
@@
where $L$ is the length of the side and $(x,y)$ is the coordinate of lattice sites. For $p=4$, $(\tilde{x},\tilde{y})$ is defined through $(x,y)=\tilde{x}\bm{e}_{1}+\tilde{y}\bm{e}_{2}$ with $\bm{e}_{1,2}$ being unit vectors along sides of the diamond [Fig. 1(a)].
We now turn to the case without spin U(1) symmetry, where the operator $\mathcal{C}$ is no longer the chiral symmetry of $H$. Therefore, $\mathcal{N}$ in Eq. (5) and $\bm{\chi}_{p}$ in Eq. (7) are no longer well-defined. However, $\bar{\mathcal{N}}$ is still well-defined if the spin mixing term is not too strong, of which the exact mathematical condition is derived in the Supplemental Material [63]. Without the spin U(1) symmetry, MKP has a $Z_{2}$ topological classification. We redefine the correspondence between the spin Bott indices and the pattern of MKP as
@@
Here, ${\bar{\mathcal{N}}}_{p}^{(i)}$ is still generated by the same $f_{p}^{(i)}$ through Eq. (6) and each element of $\bar{\boldsymbol{\chi}}_{p}$ is 0 or 2, which counts the number of topologically robust Majorana corner states at a given corner.
Model.— As an illustration, we study the Kane-Mele model with sublattice-dependent superconducting pairing potentials. The model Hamiltonian is
@@
The normal state of $\mathcal{H}$ realizes quantum spin Hall insulator when the chemical potential $\mu$ is in the bulk gap and hosts gapless helical states along edges. For simplicity, we take $\mu=0$ to analyze the edge states [63]. The helical edge states along the zigzag and bearded edges (type-I), formed by either the A or B atoms, are gapped by the superconducting pairing with magnitude $\Delta_{\text{A}}$ or $\Delta_{\text{B}}$ [63]. In comparison, the helical edge states along the armchair edge (type-II), formed by both the A and B atoms, are gapped by the superconducting potentials with magnitude $(\Delta_{\text{A}}+\Delta_{\text{B}})/2$ [63]. MKP emerge at a given corner when the pairing gaps of two adjacent edges have a sign change. This scenario can occur in two special cases, (i) both adjacent edges belong to type-I edge and are formed by A and B atoms, respectively, with $\Delta_{\text{A}}\Delta_{\text{B}}<0$; (ii) two adjacent edges belong to type-I and type-II edges, respectively, satisfying $(\Delta_{\text{A}}+\Delta_{\text{B}})\Delta_{\alpha}<0$ (type-I edge is formed by the $\alpha$ atom). For case (i) [(ii)], we find that MKP can be engineered in systems with diamond and hexagon shapes (for square and dodecagon shapes, see [63]). In the following, we focus on the case (i), and case (ii) is discussed in [63].
In the diamond- [Fig. 1] and hexagon-shaped [Fig. 2] systems, all edges belong to type-I and are formed by A or B atom, depending on the specific boundary termination. Therefore, all possible distribution patterns of MKP can be realized by varying edge cleavage to design sign patterns of edge-state energy gap. Specifically, the number of corners that host MKP can be any even number not exceeding 4 (6) for the diamond (hexagon) system. In Figs. 1(b)-1(d) and Figs. 2(a)-2(e), we present several typical spatial distribution patterns of MKP, demonstrating the flexibility to engineer MKP. We employ the vectors $\bm{\chi}_{4}$ and $\bm{\chi}_{6}$ to describe these configurations of MKP, which can be fully characterized by the vectors $\bm{\mathcal{N}}_{4}=({\mathcal{N}}_{4}^{(1)},{\mathcal{N}}_{4}^{(2)},{\mathcal{N}}_{4}^{(3)})$ and $\bm{\mathcal{N}}_{6}=({\mathcal{N}}_{6}^{(1)},{\mathcal{N}}_{6}^{(2)},{\mathcal{N}}_{6}^{(3)},{\mathcal{N}}_{6}^{(4)},{\mathcal{N}}_{6}^{(5)})$, respectively. Our numerical results are consistent with Eq. (8). In Figs. 1(e)-1(g), we present the evolution of $\bm{\mathcal{N}}_{4}$ and energies as functions of $\Delta_{\text{B}}$ for the systems depicted in Figs. 1(b)-1(d), which consistently characterizes the topological phase transitions. We emphasize that this correspondence captured by $\bm{\mathcal{N}}_{4,6}$ reflects the nontrivial topology of both bulk and edge states [60]. In contrast, the Bott index defined under the periodic boundary conditions [47] can not precisely characterize MKP as it fails to capture termination-dependent edge topology.
@@
We finally consider the $s_{z}$ symmetry breaking case with $\lambda_{\text{R}}\neq 0$. In this case, we characterize MKP by $\bar{\mathcal{N}}$ defined in Eq. (6). For instance, we consider nonzero $\lambda_{\text{R}}$ in the system associated with Fig. 1(b). In Fig. 3(b), we present the numerical results of $\bar{{\mathcal{N}}}_{4}^{(1,2,3)}$ and energies as functions of $\lambda_{v}$ at a fixed $\lambda_{\text{R}}$. The topological invariants correctly characterize the topological phase transition driven by the increasing of $\lambda_{v}$.

### v3

# Characterization of higher-order topological superconductors using Bott indices

@@ Abstract
The abundance of bulk and boundary topologies in higher-order topological phases offer remarkable tunability and diversity to boundary states but also pose a challenge to their unified topological characterization. In this work, we propose a theoretical framework to characterize time-reversal invariant topological superconductors hosting Majorana Kramers pairs (MKP) of corner states by using a series of spin Bott indices, which capture both bulk and boundary states topology. The developed invariants can characterize MKP in arbitrarily shaped systems and all distinct spatial distribution patterns of MKP. As an illustrative example, we apply our theory to analyze the Kane-Mele model with sublattice-dependent superconducting pairing potentials. In this representative model, both intrinsic and extrinsic higher-order topological superconductors can be realized and various patterns of MKP can be engineered through edge cleavage. Despite their high sensitivity to boundary terminations, MKP can be faithfully characterized by the proposed topological invariants. We further demonstrate the characterization of higher-order topological superconductors in the BDI symmetry class using Bott indices without resolving the spin degree of freedom.
I Introduction
@@
Higher-order topological phases, hosting gapless boundary states with codimension greater than one, have attracted extensive research interest in both theory [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16] and experiments [17, 18, 19, 20, 21, 22, 23]. Higher-order topological superconductors (HOTSCs) featuring Majorana corner states [24, 25, 26, 27, 28] are of particular interest because of their potential applications in topological quantum computation [29, 30, 31, 32]. Various systems have been proposed to realize HOTSCs, such as superconducting-proximitized topological insulators [33, 34, 35, 36, 37], odd-parity superconductors [38, 39, 40, 41], and iron-based superconductors [42, 43, 44, 45, 46, 47]. In contrast to first-order topology, the topological boundary states of HOTSCs typically depend on not only bulk states topology but also the boundary terminations [4]. The abundant bulk and boundary topologies of HOTSCs endow boundary states with high tunability and richness but also present the challenge of their unified topological characterizations [48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. Several theoretical frameworks for characterizing higher-order topology have been developed, such as multipole moment [58, 59, 60], chiral multipole number [50], and symmetry indicators [61, 62, 63, 64]. Despite these seminal works, a unified topological characterization of Majorana corner states remains elusive. A key difficulty is that the previously proposed invariants are defined by only bulk states while Majorana corner states are sensitive to both bulk and boundary topologies [4, 65], which hinders establishing an exact correspondence between the proposed invariants and Majorana corner states. As such, a comprehensive topological characterization of HOTSCs must encompass both bulk and boundary state topologies, a challenge that existing frameworks have yet to fully address. Moreover, no current theory provides a complete characterization of HOTSCs with arbitrary shapes or captures the diverse real-space patterns of Majorana corner states.
Bott index, deeply rooted in $K$-theoretic invariants [66, 67, 68, 69], stands as a potent instrument and has been used to characterize various topological phases of matter [70, 71, 72, 73, 74, 75, 76]. These include topological insulators characterized by winding number [66, 75], Chern insulators [70, 77], and quantum spin Hall insulators [72]. As a real space topological invariant, Bott index is defined by a couple of unitary matrices. When the unitary matrix is generated by the real space polynomial $e^{i2\pi x/L}$ and/or $e^{i2\pi y/L}$, the Bott indices can provide equivalent expressions for winding numbers or Chern numbers [75, 77]. Recently, the Bott index was used to characterize chiral symmetric higher-order topological insulators by generalizing the form of real space polynomial $e^{i2\pi f(\bm{r})}$ [50, 78]. This break through motivates us to explore the general topological characterization of HOTSCs by Bott index.
@@
In this work, we develop a theoretical framework to characterize time-reversal invariant HOTSCs by higher-order topological invariants based on a series of spin Bott indices. Our approach is motivated by recent theoretical progresses of characterization chiral symmetric systems by using Bott indices [50, 78]. Time-reversal invariant superconductors are in the DIII symmetry class and naturally host chiral symmetry $C=-iTP$, where $T$ and $P$ are, respectively, time-reversal and particle-hole symmetry. However, a direct application of the Bott index $N$ proposed in Ref. 78 based on the chiral symmetry $C$ of superconductors can not capture the higher-order topology since $N$ vanishes as required by the $T$ symmetry (see Eqs. (2) and (3)). Therefore, we develop spin Bott indices to characterize HOTSCs, in analogy to the spin-resolved topological characterization of quantum spin Hall insulators [79]. We emphasize that the Bott indices here are obtained under open boundary conditions, which enables capturing the topology of both bulk and boundary states. Particularly, the defined spin Bott indices can characterize MKP in systems with arbitrary shapes and all distinct spatial distribution patterns of MKP. To demonstrate the effectiveness of our theory, we study the Kane-Mele model [80] with sublattice-dependent superconducting pairing potentials. In this model, MKP can be flexibly engineered by boundary cleavage, as shown in Figs. 1 and 2, and all different patterns of MKP can be characterized by the proposed invariants. In addition, we also investigate disorder effects on the HOTSCs informed by the invariants. Furthermore, by using Bott indices without resolving the spin degree of freedom, we demonstrate the characterization of HOTSCs in the BDI symmetry class.
This paper is organized as follows. In Sec. II, we develop the spin Bott index for the systems of DIII symmetry class. In Sec. III, we develop the framework for characterizing MKP by spin Bott indices for systems with or without spin conservation. In Sec. IV, we use the Kane-Mele model with the sublattice-dependent superconducting pairing potentials to examine our theory. In Sec. V, we further demonstrate the topological characterization of HOTSCs in the BDI symmetry class by Bott indices. In Sec. VI, we present a brief discussion and summary. Appendices A-G complement the main text with additional technical details.
II Spin Bott indices
The symmetry operators of superconductors in the DIII symmetry class can be represented as $T=is_{y}\mathcal{K}$, $P=\tau_{y}s_{y}\mathcal{K}$, and $C=\tau_{y}$ with Pauli matrices $s_{y}$ and $\tau_{y}$ acting, respectively, on the spin and particle-hole spaces, and $\mathcal{K}$ being the complex conjugation operator. We first show that the application of the previously defined Bott indices [50, 78] to superconductors in the DIII class based on $C$ can not capture the higher-order topology.
@@ II Spin Bott indices
Using singular value decomposition $h=U_{A}\Sigma U_{B}^{\dagger}$, the Bott index can be defined in the following equivalent ways (see Appendix A for details)
@@
This restriction is consistent with the MKP related by $T$ having opposite eigenvalues of $C$. We denote the MKP at a given corner as $\gamma_{1}$ and $\gamma_{2}$, with $\gamma_{2}=T\gamma_{1}$. The Majorana corner states can always be chosen as the eigenstates of $C$ through a superposition of $\gamma_{1}$ and $\gamma_{2}$. Without loss of generality, we choose $C\gamma_{1}=\gamma_{1}$, which leads to
@@
Here $M_{z}$ commutes with $T$. With $s_{z}$ conservation, $\mathcal{N}=\bar{\mathcal{N}}$ (see Appendix B for details).
III Characterization of MKP
When $[s_{z},H]=0$, $H$ is block-diagonal in the spin space and each block belongs to the AIII symmetry class [81], allowing one corner to have $Z$ pairs of MKP. To describe the pattern of MKP at corners of a system with open-boundary conditions, we define the vector
@@ III Characterization of MKP
${f}_{4}^{(1)}=2\tilde{x}\tilde{y}/L^{2},\quad{f}_{4}^{(2)}=\tilde{x}/L,\quad{f}_{4}^{(3)}=\tilde{y}/L,$
@@
where $L$ is the length of the side and $(x,y)$ is the coordinate of lattice sites. For $p=4$, $(\tilde{x},\tilde{y})$ is defined through $(x,y)=\tilde{x}\bm{e}_{1}+\tilde{y}\bm{e}_{2}$ with $\bm{e}_{1,2}$ being unit vectors along sides of the diamond [Fig. 1(a)]. It is noted that the choice of polynomials $f_{p}^{(i)}(\bm{r})$ is not unique for certain shaped system. In Appendix C, we present an alternative choice of $f_{6}^{(i)}(\bm{r})$ with $i=1,2,3,4,5$.
We now turn to the case without spin U(1) symmetry, where the operator $\mathcal{C}$ is no longer the chiral symmetry of $H$. Therefore, $\mathcal{N}$ in Eq. (5) and $\bm{\chi}_{p}$ in Eq. (7) are no longer well-defined. However, $\bar{\mathcal{N}}$ is still well-defined if the spin mixing term is not too strong, of which the exact mathematical condition is derived in Appendix D. Without the spin U(1) symmetry, MKP has a $Z_{2}$ topological classification. We redefine the correspondence between the spin Bott indices and the pattern of MKP as
@@
Here ${\bar{\mathcal{N}}}_{p}^{(i)}$ is still generated by the same $f_{p}^{(i)}$ through Eq. (6) and each element of $\bar{\boldsymbol{\chi}}_{p}$ is 0 or 2, which counts the number of topologically robust Majorana corner states at a given corner.
IV Theoretical model
As an illustration, we study the Kane-Mele model with sublattice-dependent superconducting pairing potentials. The model Hamiltonian is
@@ IV Theoretical model
The normal state of $\mathcal{H}$ realizes quantum spin Hall insulator when the chemical potential $\mu$ is in the bulk gap and hosts gapless helical states along edges. For simplicity, we take $\mu=0$ to analyze the edge states (see Appendix E for details). The helical edge states along the zigzag and bearded edges (type-I), formed by either the A or B atoms, are gapped by the superconducting pairing with magnitude $\Delta_{\text{A}}$ or $\Delta_{\text{B}}$ (see Appendix E for details). In comparison, the helical edge states along the armchair edge (type-II), formed by both the A and B atoms, are gapped by the superconducting potentials with magnitude $(\Delta_{\text{A}}+\Delta_{\text{B}})/2$. MKP emerge at a given corner when the pairing gaps of two adjacent edges have a sign change. This scenario can occur in two special cases, (i) both adjacent edges belong to type-I edge and are formed by A and B atoms, respectively, with $\Delta_{\text{A}}\Delta_{\text{B}}<0$; (ii) two adjacent edges belong to type-I and type-II edges, respectively, satisfying $(\Delta_{\text{A}}+\Delta_{\text{B}})\Delta_{\alpha}<0$ (type-I edge is formed by the $\alpha$ atom). For case (i) [(ii)], we find that MKP can be engineered in systems with diamond and hexagon shapes (for square and dodecagon shapes). In the following, we focus on the case (i), and case (ii) is discussed in Appendix F.
@@
In the diamond- [Fig. 1] and hexagon-shaped [Fig. 2] systems, all edges belong to type-I and are formed by A or B atom, depending on the specific boundary termination. Therefore, all possible distribution patterns of MKP can be realized by varying edge cleavage to design sign patterns of edge-state energy gap. Specifically, the number of corners that host MKP can be any even number not exceeding 4 (6) for the diamond (hexagon) system. In Figs. 1(b)-1(d) and Figs. 2(a)-2(e), we present several typical spatial distribution patterns of MKP, demonstrating the flexibility to engineer MKP. We employ the vectors $\bm{\chi}_{4}$ and $\bm{\chi}_{6}$ to describe these configurations of MKP, which can be fully characterized by the vectors $\bm{\mathcal{N}}_{4}=({\mathcal{N}}_{4}^{(1)},{\mathcal{N}}_{4}^{(2)},{\mathcal{N}}_{4}^{(3)})$ and $\bm{\mathcal{N}}_{6}=({\mathcal{N}}_{6}^{(1)},{\mathcal{N}}_{6}^{(2)},{\mathcal{N}}_{6}^{(3)},{\mathcal{N}}_{6}^{(4)},{\mathcal{N}}_{6}^{(5)})$, respectively. Our numerical results are consistent with Eq. (8). In Figs. 1(e)-1(g), we present the evolution of $\bm{\mathcal{N}}_{4}$ and energies as functions of $\Delta_{\text{B}}$ for the systems depicted in Figs. 1(b)-1(d), which consistently characterizes the topological phase transitions. We emphasize that this correspondence captured by $\bm{\mathcal{N}}_{4,6}$ reflects the nontrivial topology of both bulk and edge states [78]. In contrast, the Bott index defined under the periodic boundary conditions [50] can not precisely characterize MKP as it does not capture termination-dependent edge topology.
@@
We further consider the $s_{z}$ symmetry breaking case with $\lambda_{\text{R}}\neq 0$. In this case, we characterize MKP by $\bar{\mathcal{N}}$ defined in Eq. (6). For instance, we consider nonzero $\lambda_{\text{R}}$ in the system associated with Fig. 1(b). In Fig. 3(b), we present the numerical results of $\bar{{\mathcal{N}}}_{4}^{(1,2,3)}$ and energies as functions of $\lambda_{v}$ at a fixed $\lambda_{\text{R}}$. The topological invariants correctly characterize the topological phase transition driven by the increasing of $\lambda_{v}$.
Figure 4: Majorana corner states in different-shaped systems that belong to the BDI symmetry class. The letters $\text{A}_{\text{AB}}$, $\text{Z}_{\alpha}$, $\text{B}_{\alpha}$ label the armchair, zigzag, and bearded edges, respectively, with $\alpha\in$A, B}. The insets plot the energies close to zero. In (a), we take $\Delta=0.25,\eta_{\text{B}}=0.4,\eta_{\text{A}}=0$. In (b), we take $\Delta=0.25,\eta_{\text{B}}=\eta_{\text{A}}=0.4$.
V HOTSC in the BDI symmetry class
We now turn to another important superconducting system in the BDI symmetry class, which hosts the chiral symmetry as a combination of effective time-reversal symmetry $\tilde{T}$ and particle-hole symmetry $P$ with $\tilde{T}^{2}=1$ and $P^{2}=1$. For the BDI symmety class, the Majorana corner states are protected by chiral symmetry and have a $Z$ topological classification. Therefore, the Majorana corner states can be directly characterized by the Bott index $N$ defined by Eq. (2) without the need of resolving the spin degree of freedom. The different Majorana patterns in the real space space described by $\bm{\chi}_{p}$ ($\bm{\chi}_{p}\in\bm{Z}$) can be characterized by $\bm{N}_{p}$ ($\bm{N}_{p}\in\bm{Z}$) as
$\boldsymbol{\chi}_{p}=\mathcal{M}^{-1}\cdot\left({\bm{N}}_{p},0\right)^{\mathrm{T}}.$ (12)
Here $\bm{N}_{p}=({{N}}_{p}^{(1)},\dots,{{N}}_{p}^{(p-1)})$ and ${N}_{p}^{(i)}$ is generated by the polynomial ${f}_{p}^{(i)}$ of position operator for $11\leq i\leq p-$ 1.
We note that the BDI symmetry class can be realized in two special cases by applying a magnetic Zeeman field or considering magnetic orders on top of the DIII symmetry class systems which respect spin SU$(2)$ [case (a)] or U$(1)$ [case (b)] symmetry. For the case (a), the magnetic field can be applied along arbitrary $\bm{n}$ direction and the $\tilde{T}$ symmetry is defined as $\tilde{T}=Ts_{\bm{e}}$ with $\{s_{\bm{n}},s_{\bm{e}}\}=0$. For the case (b), the magnetic field direction is constrained to be in the plane perpendicular to the spin U$(1)$ $z$-axis and $\tilde{T}$ symmetry is defined as $Ts_{z}$. The case (a) can occur in the superconducting systems with collinear magnetic structures described by spin point group [88]. Even for the coplanar magnetic structures ($xy$ plane assumed) systems, the BDI symmetry class can still be realized with the definition of $\tilde{T}=Ts_{z}$. The case (b) can be exemplified by our model (Eq. (11)) by adding an in-plane magnetic Zeeman field and setting $\lambda_{\text{R}}=0$. In the following, we focus on the HOTSCs in the BDI symmetry class for this model.
To engineer Majorana corner states, we consider site-dependent Zeeman field which can be obtained in the Kane-Mele-Hubbard model [86]. The total model Hamiltonian in momentum space can be written as
${H}(\bm{k})$ $=f_{x}(\bm{k})\tau_{z}\sigma_{x}s_{0}+f_{y}(\bm{k})\tau_{z}\sigma_{y}s_{0}+f_{z}(\bm{k})\tau_{z}\sigma_{z}s_{z}$ (13)
$+\Delta\tau_{x}s_{0}+\eta_{\text{A}}\tau_{0}(\sigma_{0}+\sigma_{z})/2s_{x}+\eta_{\text{B}}\tau_{0}(\sigma_{0}-\sigma_{z})/2s_{x},$
where $f_{x,y,z}$ are momentum-dependent functions (see Appendix E) and Pauli matrices $\sigma$ act on the orbitals. We assume an $s$-wave pairing with $\Delta_{\text{A}}=\Delta_{\text{B}}=\Delta$. $\eta_{\text{A}}$ and $\eta_{\text{B}}$ denote the magnitude of Zeeman field at the A and B sites, respectively. $H$ respects the $\tilde{T}$ and $P$ symmetries, where $\tilde{T}=Ts_{z}=-s_{x}K$ and $P=\tau_{y}s_{y}K$. The chiral symmetry can be represented by combination of $\tilde{T}$ and $P$ symmetries, $\mathcal{C}=i\tilde{T}P=\tau_{y}s_{z}$. Thus, $H$ in Eq. (13) belongs to the BDI symmetry class. In the following, we show that various distinct spatial patterns of Majorana corner states can be realized by tuning $\Delta,\eta_{\text{A}}$ and $\eta_{\text{B}}$.
According to the edge theory analysis (see Appendix E), the gapless edge states at the Dirac points along the armchair and zigzag directions are the approximate eigenstates of operators $\sigma_{y}s_{z}$ and $\sigma_{z}$, respectively. Thus, the edge states along the armchair direction can be gapped by the terms $\Delta\tau_{x}s_{0},\eta_{\text{A}}\sigma_{z}s_{x}/2$, and $-\eta_{\text{B}}\sigma_{z}s_{x}/2$, which give rise to the edge states energy gap $\Delta-|\eta_{\text{A}}-\eta_{\text{B}}|/2$. The edge states along the zigzag (beard) direction formed by the $\alpha$ atom can be gapped by the terms $\Delta\tau_{x}s_{0}$ and $\eta_{{\alpha}}(\sigma_{0}+\xi\sigma_{z})s_{x}/2$ with $\xi=1,-1$ for $\alpha=\text{A},\text{B}$, which give rise to the edge states energy gap $\eta_{\alpha}-\Delta$. When $\eta_{\text{A}}=\eta_{\text{B}}=\eta$, our analysis reproduces the results obtained in Ref. 35.
The Majorana corner states can be realized when the two adjacent edges have a Zeeman-dominated and superconductivity-dominated gap, respectively. This scenario can be realized in two special cases, (i) both adjacent edges belong to type-I edge and are formed by A and B atoms, respectively, with $(|\eta_{\text{A}}|-|\Delta|)(|\eta_{\text{B}}|-|\Delta|)<0$; (ii) two adjacent edges belong to type-I and type-II edges, respectively, satisfying $|\eta_{\alpha}|>|\Delta|>|\eta_{\text{A}}-\eta_{\text{B}}|/2$ (type-I edge is formed by the $\alpha$ atom). Thus, by tuning $\Delta,\eta_{\text{A}}$ and $\eta_{\text{B}}$, various distinct spatial patterns of the Majorana corner states can be realized in the diamond, square, hexagon, and dodecagon systems. In Fig.  4, we present certain Majorana patterns in the diamond and square systems. These different Majorana patterns can be characterized by the corresponding $\bm{N}_{4}$ vectors, as listed in Fig.  4.
VI Discussion and summary
We discuss some details of our theoretical approach. Although our numerical results are given for system with certain geometry, our theory is applicable for arbitrarily shaped system. For a general system with $p$ corners, we can construct $(p-1)$ polynomials $f_{p}^{i}(\boldsymbol{r})$ that satisfy the required conditions (see the work [78] for details). Based on the exact correspondence between the generated Bott indices and patterns of Majorana corner modes, we can fully characterize higher-order topological superconductors.
We emphasize that crystalline symmetries can generally impose constraints on the Bott indices by restricting the patterns of Majorana corner modes that to be compatible with the symmetry. For example, in a time-reversal invariant system with four-fold rotation symmetry, MKP patterns that violate this symmetry are prohibited. Specifically, MKP must appear simultaneously at all four corners or not at all. Therefore, the constraints on the spin Bott index yield $\mathcal{N}_{4}^{(2)}=\mathcal{N}_{4}^{(3)}=0$ and $\mathcal{N}_{4}^{(1)}=4Z$ with $Z$ being an integer. This principle extends to other crystalline symmetries and constrains the possible values of the Bott indices. A detailed exploration of specific constraints for various crystalline symmetries is an interesting direction for future work.